[figure: 9.7 cm]

SUPERCONDUCTING CRITICAL TEMPERATURE OF LAYERED MERCUROCUPRATE FAMILY

A.L. Kuzemsky^{a}, I.G. Kuzemskaya^{b} and A.A. Cheglokov^{a} aBogolubov Laboratory of Theoretical Physics, Joint Institute for Nuclear Research 141980 Dubna, Moscow Region, Russia bHigh Pressure Physics Institute, Russian Academy of Sciences, 142092 Troitsk, Moscow Region, Russia

We study the physics of the mercurocuprate family from the structural point of view. We present a phenomenological approach determining the critical temperatures for homologous series HgBa_{2}Ca_{n−1}Cu_{n}O_{2n+2+δ} depending of number of layers n. The model is based on the Lawrence-Doniach theory of layered superconductors modified for the case of inequivalent layers. The redistribution of charge was taken into account. This lead to observable nonmonotonic ”bell”-shaped dependence of T_{c}(n) and provides a quantitative explanation of the experiments.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Dr. A.L. Kuzemsky

Address for correspondence : Bogolubov Laboratory of Theoretical Physics Joint Institute for Nuclear Research 141980 DUBNA, Moscow Region, Russia

Email address : kuzemsky@thsun1.jinr.ru